Lean Accounting Dashboard Inc. manufactures and assembles automobile instrument panels for both eCar Motors and Greenville Mo1. Determine the budgeted cell conversion cost per hour. per hour 2. Determine the budgeted cell conversion cost per unit. $4. Determine the ending balance in Raw and In Process Inventory and Finished Goods Inventory. Raw and In Process Inventory: FFor #3, the Labels are: Accounts Payable, Accounts Receivable, Conversion Costs, Cost of Goods Sold, Finished Goods Inventory, Raw and In-Process Inventory, Raw Materials, Fixed Assets, Sales, Sales Returns and Allowances.
Answer #1
1
Budgeted cell conversion cost per hour 700 per hour =1400000/2000
2
Budgeted cell conversion cost per unit 280.00 per unit =700*(24/60)
3
Raw and in process inventory 108000 =450*240
     Accounts Payable 108000
Raw and in process inventory 112000 =400*280
       Conversion costs 112000
Finished goods inventory 197600 =380*(240+280)
       Raw and in process inventory 197600
Accounts Receivable 280000 =350*800
       Sales 280000
Cost of goods sold 182000 =350*(240+280)
        Finished goods inventory 182000
4
Raw and In-Process Inventory 22400 =108000+112000-197600
Finished goods inventory 15600 =197600-182000
5
Lean accounting is different from traditional accounting because it is more simplified and uses minimal control. As a result, the number of transactions are reduced . In many lean operations, purchased materials are charged to a “raw and in process inventory” account . Direct labor is frequently included as a conversion cost of the cell . Often, nonfinancial performance measures, such as lead time or quality measures , are used to monitor performance.